Circular No. 6096

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

EXO 1846–031 S. N. Zhang, B. A. Harmon, C. A. Wilson, G. J. Fishman, and W. S.

Paciesas, Marshall Space Flight Center, NASA; J. E. Grindlay and D. Barret, Harvard-Smithsonian Center for Astrophysics (CfA); and E. Ford, M. Tavani, and P. Kaaret, Columbia University, report: “A recent outburst from the x-ray transient EXO 1846–031 has been detected by BATSE. An earth-occultation image, integrated between Sept. 21 and 30, showed a significant peak consistent with the location of EXO 1846–031. The location uncertainty is 0^{o}.1 perpendicular to a line between the source location (α = 18^{h}49^{m}.3, δ = −3^{o}04′, equinox 2000.0) and α = 18^{h}56^{m}, δ = +6^{o}.0, and extends ∼ ±1^{o} along the line centered on the source. No other catalogued hard x-ray sources are consistent with the bright peak. Although the statistical accuracy is poor, the outburst began around Sept. 15 ± 3 and peaked on Sept. 23 ± 3. As of Oct. 8 ± 3 days, the source has dropped below the sensitivity limit of BATSE. This source is a black-hole candidate because of its ultrasoft spectral component and hard x-ray tail. Between Sept. 21 and Oct. 1, the spectrum could be acceptably fitted by a power law at least to 100 keV with an index of −2.9 ± 0.2. The average flux (band 20–100 keV) during this interval was ∼ 80 ± 20 mCrab. We note that the original EXOSAT observations (Parmar et al. 1993, A.Ap. 279, 179) showed a decay constant for soft x-rays of 85 days. The BATSE power law is consistent with the 1985 EXOSAT observation of a hard tail above 10 keV. Immediate follow-up observations at optical (the counterpart has not been identified yet), radio, and soft x-ray wavelengths are strongly encouraged.”

J. Grindlay, M. Garcia, and P. Zhao, CfA, report: “B. McLean obtained moderately deep I-band imaging of the EXOSAT error circle for EXO 1846–031 (Parmar et al., ibid.) with the 1.2-m telescope at the Whipple Observatory on Oct. 14.15 UT. No optical counterpart brighter than I = 21.5 in the EXOSAT error circle was detected, suggesting that the source is either very distant or heavily absorbed or that this outburst decayed more quickly. Deeper observations would still be useful.”

COMET MACHHOLZ (1994r) Total visual magnitude estimate by C. S. Morris, Pine Mountain Club,

CA (0.26-m reflector): Oct. 12.47 UT, 10.6 (5′ coma).

1994 October 15 Daniel W. E. Green